# Yutong Xia [Uncaptioned image] [Uncaptioned image]  Ao Qu [Uncaptioned image]  Yunhan Zheng [Uncaptioned image]  Yihong Tang [Uncaptioned image]  Dingyi Zhuang [Uncaptioned image] Yuxuan Liang [Uncaptioned image]  Cathy Wu [Uncaptioned image]  Roger Zimmermann [Uncaptioned image]  Jinhua Zhao[Uncaptioned image]

### v1

@@ Abstract
Yutong Xia [Uncaptioned image] [Uncaptioned image]  Ao Qu [Uncaptioned image]  Yunhan Zheng [Uncaptioned image]  Yihong Tang [Uncaptioned image]  Dingyi Zhuang [Uncaptioned image] Yuxuan Liang [Uncaptioned image]  Cathy Wu [Uncaptioned image]  Roger Zimmermann [Uncaptioned image]  Jinhua Zhao[Uncaptioned image] [Uncaptioned image] MIT  [Uncaptioned image] NUS  [Uncaptioned image] McGill  [Uncaptioned image] HKUST (GZ) {yutongx,qua,yunhan,dingyi,cathywu,jinhua}@mit.eduyihong.tang@mail.mcgill.ca yuxliang@outlook.comdcsrz@nus.edu.sg
@@ 1 Introduction
Recent advances in Large Language Models (LLMs) Achiam et al. (2023); Touvron et al. (2023); Team et al. (2023); Guo et al. (2025) and Multimodal Large Language Models (MLLMs) Wu et al. (2023) offer a compelling opportunity. LLMs demonstrate strong capabilities in understanding and generating text Radford et al. (2018), reasoning Wei et al. (2022); Jaech et al. (2024), planning Huang et al. (2022), code execution Chen et al. (2021), and tool use Schick et al. (2023); Qin et al. (2024). MLLMs extend these functions across diverse data modalities such as imagery Liu et al. (2023), audio Chu et al. (2023), graph Tang et al. (2024), and structured charts Han et al. (2023), lowering the barriers to engaging with complex urban data at scale Liang et al. (2025). More recently, through automated hypothesis generation, integrated analysis of multimodal data, and adaptive causal study design, LLMs enable a shift from narrow, discipline-bound efforts to more inclusive and exploratory research Lu et al. (2024); Schmidgall et al. (2025); Gottweis et al. (2025).
@@ 2 Landscape of Causal Inference in Urban Studies
This section provides a quantitative overview of the field based on a comprehensive empirical analysis. It identifies key trends in causal inference research across urban topics, data usage, methodological approaches, and geographic focus. These findings lay the foundation for designing AutoUrbanCI, our proposed AI framework for automated causal research, ensuring it is aligned with actual research needs, practices, and gaps.
@@ 3.3 Multi-agent system example
This section presents a concrete instantiation of AutoUrbanCI through a multi-agent system (MAS) example, illustrating how agents collaborate across the causal workflow in a realistic research scenario. We focus on a case study investigating the impact of congestion pricing on low-income commuters in New York City (NYC), as shown in Figure 5. This example demonstrates one possible path through the framework: Practice-Driven + Theory-Driven $\rightarrow$ Web-Based Retrieval $\rightarrow$ Quasi-Experimental Methods $\rightarrow$ Validity Check + Policy Narrative + Robustness Test. The system includes seven specialized agents—Urban Scientist, Reader, Data Engineer, Validator, Data Scientist, Experimenter, and Writer—each responsible for a distinct function in the pipeline. Detailed roles and responsibilities are described in Appendix C.
@@ 3.4 Challenges
The first challenge arises from the inherent complexity of urban data, which are multi-source, multi-granular, and multi-modal. Specifically, urban data originate from highly heterogeneous sources, including social media, satellite imagery, mobility logs, environmental sensors, and policy documents. These datasets differ not only in modality (e.g., text, images, time series, and tabular formats) but also in granularity, ranging from neighborhood-level attributes to real-time street-level observations Zhang et al. (2024b); Zou et al. (2025). Integrating such data presents substantial challenges in preprocessing, normalization, and semantic alignment Martínez-García and Hernández-Lemus (2022); Liang et al. (2025). Recent work has explored using LLMs to construct urban knowledge graphs from unstructured data, as in UrbanKGent Ning and Liu (2024), which proposes an agent framework based on triplet extraction and spatio-temporal reasoning. While promising, such approaches remain at an early stage and require further development for robust, large-scale deployment.
@@ 3.5 Evaluation
Method Development Clarity, Validity, Rigor, Innovativeness, Generalizability Human, LLM ResearchAgent (2024), MLR-Copilot (2024a)
@@
Experiment Execution Performance Improvement, Success Rate Prototype Code MLR-Copilot (2024a)
@@
For AutoUrbanCI, in addition to the existing evaluation methods, its inherent nature requires several additional evaluation focuses, including hypothesis generation significance, data processing quality, and causal inference validity. First, for idea generation evaluation, in addition to novelty and feasibility, it should also consider scientific rigor (i.e., ensuring asking appropriately clear and specific research questions adheres to established causal principles Bailey et al. (2024)) and interdisciplinary value (i.e., assessing whether they integrate insights from multiple domains, such as urban science, economics, and environmental studies, to provide innovative perspectives). Second, for data processing quality, its evaluation can focus on data completeness, consistency, preprocessing accuracy, scalability, and computational efficiency, ensuring robust, efficient, and effective multimodal data integration. Lastly, causal inference evaluation assesses causal structure feasibility, intervention validity, counterfactual consistency, and robustness to confounding. This can be achieved through synthetic data benchmarking, expert validation, or curated datasets derived from existing urban causal inference studies, ensuring that discovered causal insights are both statistically sound and practically relevant for urban policy applications.
@@ 4.2 New Human-AI Collaboration Paradigm
Traditionally, AI systems have served as passive tools in urban research workflows, used for predictive modeling Li et al. (2024b), spatial visualization Zhang et al. (2024a), or data preprocessing Liang et al. (2025). However, the introduction of AutoUrbanCI represents a shift toward a more collaborative paradigm. In this framework, AI agents are no longer limited to executing predefined tasks; instead, they actively participate in the scientific reasoning process. Rather than merely accelerating isolated stages of analysis, these agents engage in context-aware interactions that reflect an evolving role: from tool to collaborator. This shift invites deeper questions about how tasks and responsibilities are distributed across human–agent systems, detailed as below.
@@
Moving beyond the traditional pipeline of human-led design, AutoUrbanCI embraces a configurable division of labor between humans and agents. Recent systems Gottweis et al. (2025); Li et al. (2024a); Lu et al. (2024); Swanson et al. (2024) reveal a common design principle: human experts define goals, constraints, and evaluation criteria, while agents handle iterative generation, optimization, and evidence organization. In these systems, natural language interfaces serve as the medium for task specification, correction, and refinement, reinforcing the human’s role as an agenda-setter and critical reviewer. In AutoUrbanCI, this flexible structure allows agents to operate independently for routine tasks such as data preprocessing or method selection, while preserving human judgment for hypothesis formulation, identification strategy, and assumption checking. Multi-agent mechanisms such as self-critique and ranking Lu et al. (2024) can further enhance robustness, but the final authority remains with human users, who validate outputs and determine their scientific and policy relevance. Rather than replacing human reasoning, this structure reinforces it: agents scaffold the causal workflow, while humans remain responsible for framing questions, defining standards, and interpreting outcomes in context.
@@ B.2 LLMs & MLLMs
Multimodal Large Language Models (MLLMs) extend this paradigm by incorporating additional modalities, e.g., imagery Liu et al. (2023); Yan et al. (2024), audio Chu et al. (2023), graph Tang et al. (2024), and structured charts Han et al. (2023), alongside text. These non-text inputs are embedded into a shared representation space as $\mathbf{e}_{V}$, while the textual prompt is embedded as $\mathbf{e}_{T}$. The MLLM then generates output tokens $\mathbf{y}=(y_{1},y_{2},\dots,y_{L})$ according to:
@@
Appendix C Roles and Functions of Agents in the Example Use Case

### v2

# Yutong Xia[Uncaptioned image][Uncaptioned image]  Ao Qu[Uncaptioned image]  Yunhan Zheng[Uncaptioned image]  Yihong Tang[Uncaptioned image]  Dingyi Zhuang[Uncaptioned image] Yuxuan Liang[Uncaptioned image]  Shenhao Wang[Uncaptioned image]  Cathy Wu[Uncaptioned image]  Lijun Sun[Uncaptioned image] Roger Zimmermann[Uncaptioned image]  Jinhua Zhao[Uncaptioned image]

@@ Abstract
Yutong Xia[Uncaptioned image][Uncaptioned image]  Ao Qu[Uncaptioned image]  Yunhan Zheng[Uncaptioned image]  Yihong Tang[Uncaptioned image]  Dingyi Zhuang[Uncaptioned image] Yuxuan Liang[Uncaptioned image]  Shenhao Wang[Uncaptioned image]  Cathy Wu[Uncaptioned image]  Lijun Sun[Uncaptioned image] Roger Zimmermann[Uncaptioned image]  Jinhua Zhao[Uncaptioned image] [Uncaptioned image]MIT  [Uncaptioned image]NUS  [Uncaptioned image]McGill  [Uncaptioned image]HKUST (GZ)  [Uncaptioned image]UFL {yutongx,qua,yunhan,dingyi,cathywu,jinhua}@mit.eduyihong.tang@mail.mcgill.ca lijun.sun@mcgill.cayuxliang@outlook.comshenhaowang@ufl.edudcsrz@nus.edu.sg
@@ 1 Introduction
Recent advances in Large Language Models (LLMs) Achiam et al. (2023); Touvron et al. (2023); Team et al. (2023); Guo et al. (2025) and Multimodal Large Language Models (MLLMs) Wu et al. (2023) offer a compelling opportunity. LLMs demonstrate strong capabilities in understanding and generating text Radford et al. (2018), reasoning Wei et al. (2022); Jaech et al. (2024), planning Huang et al. (2022), code execution Chen et al. (2021), and tool use Schick et al. (2023); Qin et al. (2024). MLLMs extend these functions across diverse data modalities such as imagery Liu et al. (2023a), audio Chu et al. (2023), graph Tang et al. (2024), structured charts Han et al. (2023), and unstructured reports Li et al. (2024a), lowering the barriers to engaging with complex urban data at scale Liang et al. (2025). More recently, through automated hypothesis generation, integrated analysis of multimodal data, and adaptive causal study design, LLMs enable a shift from narrow, discipline-bound efforts to more inclusive and exploratory research Lu et al. (2024); Schmidgall et al. (2025); Gottweis et al. (2025).
@@ 2 Landscape of Causal Inference in Urban Studies
This section provides a quantitative overview of the field based on a comprehensive empirical analysis. It identifies key trends in causal inference research across urban topics, data usage, methodological approaches, and geographic focus. These findings lay the foundation for designing AutoUrbanCI, our proposed AI framework for automated causal research, ensuring it is aligned with actual research needs, practices, and gaps (see Appendix C for methodological details).
@@ 3.3 Multi-agent system example
This section presents a concrete instantiation of AutoUrbanCI through a multi-agent system (MAS) example, illustrating how agents collaborate across the causal workflow in a realistic research scenario. We focus on a case study investigating the impact of congestion pricing on low-income commuters in New York City (NYC), as shown in Figure 5. This example demonstrates one possible path through the framework: Practice-Driven + Theory-Driven $\rightarrow$ Web-Based Retrieval $\rightarrow$ Quasi-Experimental Methods $\rightarrow$ Validity Check + Policy Narrative + Robustness Test. The system includes seven specialized agents—Urban Scientist, Reader, Data Engineer, Validator, Data Scientist, Experimenter, and Writer—each responsible for a distinct function in the pipeline. Detailed roles and responsibilities are described in Appendix D.
@@ 3.4 Challenges
The first challenge arises from the inherent complexity of urban data, which are multi-source, multi-granular, and multi-modal. Specifically, urban data originate from highly heterogeneous sources, including social media, satellite imagery, mobility logs, environmental sensors, and policy documents. These datasets differ not only in modality (e.g., text, images, time series, and tabular formats) but also in granularity, ranging from neighborhood-level attributes to real-time street-level observations Zhang et al. (2024b); Zou et al. (2025). Integrating such data presents substantial challenges in preprocessing, normalization, and semantic alignment Martínez-García and Hernández-Lemus (2022); Liang et al. (2025). Recent work has explored construct urban knowledge graphs from unstructured data Liu et al. (2023b); Ning and Liu (2024), which aim to represent complex urban entities and their relationships in a structured, machine-readable form. While promising, such efforts differ from our objective: rather than building relational graphs, our framework seeks to convert multi-modal observations into structured, temporally aligned datasets that support causal inference experiments.
@@ 3.5 Evaluation
Method Development Clarity, Validity, Rigor, Innovativeness, Generalizability Human, LLM ResearchAgent (2024), MLR-Copilot (2024b)
@@
Experiment Execution Performance Improvement, Success Rate Prototype Code MLR-Copilot (2024b)
@@
For AutoUrbanCI, in addition to the existing evaluation methods, its inherent nature requires several additional evaluation focuses. We highlight three critical aspects: hypothesis generation significance, data processing quality, experiment design and causal inference validity. First, for idea generation evaluation, in addition to novelty and feasibility, it should also consider scientific rigor (i.e., ensuring asking appropriately clear and specific research questions adheres to established causal principles Bailey et al. (2024)) and interdisciplinary value (i.e., assessing whether they integrate insights from multiple domains, such as urban science, economics, and environmental studies, to provide innovative perspectives). Second, for data processing quality, its evaluation can focus on data completeness, consistency, preprocessing accuracy, scalability, and computational efficiency, ensuring robust, efficient, and effective multimodal data integration. Third, experiment design requires assessing the system’s ability to select appropriate causal identification strategies (e.g., DiD, IV), justify assumptions, and choose valid baselines. One practical evaluation approach is to reproduce classic urban causal inference studies and compare whether the agent’s design choices align with or diverge from established methodological best practices. Lastly, causal inference evaluation assesses causal structure feasibility, intervention validity, counterfactual consistency, and robustness to confounding. This can be achieved through synthetic data benchmarking, expert validation, or curated datasets derived from existing urban causal inference studies, ensuring that discovered causal insights are both statistically sound and practically relevant for urban policy applications. Table 2 summarizes the full set of evaluation dimensions and highlights the ones that are particularly specific to the AutoUrbanCI setting.
Evaluation Focus Key Aspects Specific?
Hypothesis Quality Causal clarity, scientific rigor, cross-domain relevance
Data Processing Completeness, temporal consistency, multimodal fusion quality
Causal Inference Rigor Intervention soundness, robustness checks, counterfactual consistency
Experiment Design Appropriate method choice, identification strategy, baseline comparisons
Reproducibility & Transparency Log traceability, reproducible pipeline setup, transparent assumptions
Ethical & Social Considerations Fairness, bias detection, value alignment, responsible intervention suggestion
System Usability & Efficiency Latency, scalability, agent coordination efficiency, task success rate
Table 2: Evaluation dimensions for AutoUrbanCI. The final column marks criteria that are particularly specific to AutoUrbanCI.
@@ 4.2 New Human-AI Collaboration Paradigm
Traditionally, AI systems have served as passive tools in urban research workflows, used for predictive modeling Li et al. (2024c), spatial visualization Zhang et al. (2024a), or data preprocessing Liang et al. (2025). However, the introduction of AutoUrbanCI represents a shift toward a more collaborative paradigm. In this framework, AI agents are no longer limited to executing predefined tasks; instead, they actively participate in the scientific reasoning process. Rather than merely accelerating isolated stages of analysis, these agents engage in context-aware interactions that reflect an evolving role: from tool to collaborator. This shift invites deeper questions about how tasks and responsibilities are distributed across human–agent systems, detailed as below.
@@
Moving beyond the traditional pipeline of human-led design, AutoUrbanCI embraces a configurable division of labor between humans and agents. Recent systems Gottweis et al. (2025); Li et al. (2024b); Lu et al. (2024); Swanson et al. (2024) reveal a common design principle: human experts define goals, constraints, and evaluation criteria, while agents handle iterative generation, optimization, and evidence organization. In these systems, natural language interfaces serve as the medium for task specification, correction, and refinement, reinforcing the human’s role as an agenda-setter and critical reviewer. In AutoUrbanCI, this flexible structure allows agents to operate independently for routine tasks such as data preprocessing or method selection, while preserving human judgment for hypothesis formulation, identification strategy, and assumption checking. Multi-agent mechanisms such as self-critique and ranking Lu et al. (2024) can further enhance robustness, but the final authority remains with human users, who validate outputs and determine their scientific and policy relevance. Rather than replacing human reasoning, this structure reinforces it: agents scaffold the causal workflow, while humans remain responsible for framing questions, defining standards, and interpreting outcomes in context.
@@ B.2 LLMs & MLLMs
Multimodal Large Language Models (MLLMs) extend this paradigm by incorporating additional modalities, e.g., imagery Liu et al. (2023a); Yan et al. (2024), audio Chu et al. (2023), graph Tang et al. (2024), and structured charts Han et al. (2023), alongside text. These non-text inputs are embedded into a shared representation space as $\mathbf{e}_{V}$, while the textual prompt is embedded as $\mathbf{e}_{T}$. The MLLM then generates output tokens $\mathbf{y}=(y_{1},y_{2},\dots,y_{L})$ according to:
@@
Appendix C Methodology for Systematic Review
In Section 2, we present a comprehensive systematic review that maps the current landscape of causal inference in urban studies. Here we detail the methodology used to conduct the review as follows.
Goal.
Our objective was to obtain a quantitative understanding of the distribution of urban causal inference studies across five key dimensions: research topic, geographic focus, data modality, experiment type, and methodological approach.
Step 1: Paper Parsing via LLMs.
We developed a structured prompt to guide GPT through the analysis of each paper, with the aim of extracting relevant information and supporting evidence corresponding to each dimension. To retrieve content from full-text PDFs, we utilized the OpenAI API’s file upload feature [^5] https://platform.openai.com/docs/guides/pdf-files?api-mode=chat, which allows direct querying over document contents. This step produced initial annotations per dimension for every paper.
Step 2: Iterative Label Refinement.
For each dimension, we designed an iterative clustering pipeline that leverages GPT-4o to semantically group similar entries and propose representative labels. The algorithm identifies recurring patterns, merges equivalent concepts, and refines the terminology. This step outputs a set of standardized keywords describing each paper across the five dimensions.
Step 3: Human Validation.
To ensure accuracy and domain relevance, we manually reviewed and refined the final set of clusters and labels using expert judgment. This final validated dataset served as the basis for the visualizations and quantitative analyses presented in the main text.
Appendix D Roles and Functions of Agents in the Example Use Case